# The Geometry of Intelligence: Deterministic Functional Topology as a Foundation for Real-World Perception

### v1

@@ Abstract
Real-world phenomena do not generate arbitrary variability: their signals concentrate on compact, low-variability subsets of functional space, enabling rapid generalization from few examples. A small child can recognize a dog after only limited exposure because the perceptual manifold of “dog” is compact, structured, and low-dimensional. We formalize this principle through a deterministic functional–topological framework in which the set of valid realizations produced by a physical process forms a compact subset of a Banach space, endowed with stable invariants, a finite Hausdorff radius, and an induced continuous perceptual functional. This geometry provides explicit limits on knowledge, conditions for identifiability, and guarantees for generalization from sparse evidence—properties fundamental to both natural and artificial intelligence. Across electromechanical, electrochemical, and physiological domains, we show that real-world processes consistently generate compact perceptual manifolds with the same geometric characteristics. Their boundaries can be discovered in a fully self-supervised manner as the empirical radius saturates with increasing sampling, even when the governing equations are unknown. These results demonstrate that deterministic functional topology offers a unified mathematical foundation for perception, representation, and world-model construction. It provides a geometric explanation for why biological learners and self-supervised AI systems can generalize from few observations, and establishes compact perceptual manifolds as a fundamental building block for future AI architectures. Finally, this work unifies biological perception and modern self-supervised architectures under a single geometric principle: both derive their generalization ability from the compactness and invariants of real-world perceptual manifolds.
@@ 3.2 Universal approximation of perceptual mappings
Theorem 3.2 (Universal Approximation of Perception Functionals).
Let $\mathcal{M}$ be compact and $\Phi:\mathcal{M}\to\mathbb{R}$ be continuous. Then for every $\varepsilon>0$ there exists a neural network or other universal approximator $N_{\varepsilon}$ satisfying
$\sup_{x\in\mathcal{M}}|\Phi(x)-N_{\varepsilon}(x)|<\varepsilon.$
@@
This shows that learnability is a consequence not of architectural choices but of the underlying geometry of the perceptual manifold.
@@ 5.1 Monte Carlo radius estimation
This MC-based procedure is applied identically to real datasets and to synthetic signals generated by the simulator, enabling direct comparison of radius saturation, geometric compactness, and cross-manifold behaviour.
@@ 5.4 Practical implementation: incremental radius estimation
5.5 Cross-manifold evaluation
The Monte Carlo estimator described above provides a unified tool to quantify radius saturation and geometric compactness within a single domain. To compare different physical systems—or real versus simulated signals—we evaluate cross-manifold distances under the same cosine geometry. Small bidirectional Hausdorff distances indicate that two manifolds share the same geometric support; moderate asymmetric distances reveal partial overlap; and distances near the metric upper bound (e.g. cosine = 1) indicate that manifolds are geometrically incompatible. This procedure allows us to determine whether a simulator faithfully reproduces the admissible perceptual geometry of a real system.
@@ 6.1 Point Machines: Functional Manifold Geometry
Our analysis proceeds in three stages:
@@
  3. 3.
    cross-manifold Hausdorff distance between real and simulated manifolds.
@@ 6.1.1 Saturation of the Real PM Manifold
representing respectively: self-Hausdorff consistency, maximal internal radius, mean internal radius, and bounding-box volume.
@@ 6.1.2 Saturation of the Simulated PM Manifold
The simulator therefore produces signals lying on a compact functional manifold with the same structural constraints as real PMs.
Figure 5: Saturation of the simulated point-machine manifold. The geometry is stable and compact, mirroring the real dataset.
6.1.3 Cross-Manifold Geometry: Real vs. Simulated
We compute the bidirectional Hausdorff distances:
$d_{H}(X_{\mathrm{real}},X_{\mathrm{sim}}),\qquad d_{H}(X_{\mathrm{sim}},X_{\mathrm{real}}),$
as a function of the simulation sample size $n\in\{10,20,50,\dots,10^{4}\}$.
Both distances remain bounded and stable:
$d_{H}(\mathrm{real}\to\mathrm{sim})\in[0.03,\,0.07],\qquad d_{H}(\mathrm{sim}\to\mathrm{real})\in[0.07,\,0.11].$
Three conclusions follow:
  • The real and simulated manifolds are distinct but close. Differences correspond mainly to amplitude and noise structure.
  • The distances do not grow with $n$, proving that adding more MC samples does not deform the geometry.
  • The curves flatten completely, indicating geometric convergence.
This demonstrates that the simulator generates a manifold that is a stable Lipschitz deformation of the real one: the topology is shared, the shape is preserved, and only scaling/noise differ.
Figure 6: Cross-manifold Hausdorff distances between real and simulated PM signals. Distances are small, bounded, and saturation-stable, demonstrating that the simulation manifold is a deformation of the real one.
@@ 6.1.4 Summary of Findings
  • Cross-manifold Hausdorff distances are small and bounded.
  • MC simulation recovers the geometry of the real manifold.
@@
6.2 Batteries
Show discharge curves forming a low-dimensional compact manifold.
6.3 ECG
Show heartbeat morphology consistency and compactness.
6.4 Cross-domain consistency
Demonstrate that all domains satisfy the same geometric conditions.
@@ 9 Limitations and Future Work
Let $\Phi:\mathcal{M}\to\mathbb{R}$ be continuous and $\mathcal{M}$ be compact in $C^{0}([0,T])$.
By the Universal Approximation Theorem [cybenko1989approximation, hornik1991approximation], for any continuous function defined on a compact subset of $\mathbb{R}^{n}$, neural networks with a non-polynomial activation function can approximate it arbitrarily well.
To apply this result, observe that $\mathcal{M}$ is compact in a separable Banach space. By standard embedding arguments, continuous functionals on $\mathcal{M}$ can be approximated arbitrarily well by finite-dimensional neural approximators acting on sampled versions of the signals. Since $\Phi$ is uniformly continuous (Proposition 3.1), the approximation error remains uniformly controlled across $\mathcal{M}$.
Hence, for every $\varepsilon>0$, there exists a neural network $N_{\varepsilon}$ such that:
$\sup_{x\in\mathcal{M}}|\Phi(x)-N_{\varepsilon}(x)|<\varepsilon.$

### v2

# The Geometry of Intelligence A Functional–Topological Framework for Perception and Representation

@@ Abstract
Real-world phenomena do not generate arbitrary variability: their signals concentrate on compact, low-variability subsets of functional space, enabling rapid generalization from few examples. A small child can recognize a dog after extremely limited exposure because the perceptual manifold of “dog” is compact, structured, and low-dimensional. We formalize this principle through a deterministic functional–topological framework in which the set of valid realizations produced by a physical process forms a compact subset of a Banach space, endowed with stable invariants, a finite Hausdorff radius, and an induced continuous perceptual functional.
This geometry provides explicit limits on knowledge, conditions for identifiability, and guarantees for generalization from sparse evidence—properties fundamental to both natural and artificial intelligence. Across electromechanical, electrochemical, and physiological domains, we show that real-world processes consistently generate compact perceptual manifolds with the same geometric characteristics. Their boundaries can be discovered in a fully self-supervised manner as the empirical radius saturates with increasing sampling, even when the governing equations are unknown.
These results demonstrate that deterministic functional topology offers a unified mathematical foundation for perception, representation, and world-model construction. It provides a geometric explanation for why biological learners and self-supervised AI systems can generalize from few observations, and establishes compact perceptual manifolds as a fundamental building block for future AI architectures. Finally, this work unifies biological perception and modern self-supervised models under a single geometric principle: both derive their generalization ability from the compactness and invariants of real-world perceptual manifolds.
Keywords: Functional topology; Perceptual manifolds; Geometric foundations of intelligence; Compactness and invariants; Deterministic generative processes; Self-supervised learning geometry; Hausdorff radius; Representation learning.
@@ 3.2 Universal approximation of perceptual mappings
Theorem 3.2 (Universal Approximation on a Compact Perceptual Manifold).
Let $\mathcal{M}\subset C^{0}([0,T])$ be compact and let $\Phi:\mathcal{M}\to\mathbb{R}$ be continuous. Then for every $\varepsilon>0$ there exists a finite-dimensional embedding $\pi_{N}:\mathcal{M}\to\mathbb{R}^{N}$ and a universal approximator $N_{\varepsilon}:\mathbb{R}^{N}\to\mathbb{R}$ such that
$\sup_{x\in\mathcal{M}}|\Phi(x)-N_{\varepsilon}(\pi_{N}(x))|<\varepsilon.$
@@
This shows that learnability arises from the compact geometry of the perceptual manifold: once the domain of admissible signals is compact and perceptual mappings are continuous, finite models suffice to approximate perception arbitrarily well.
@@ 5.1 Monte Carlo radius estimation
This MC-based procedure is applied identically to real datasets and to synthetic signals generated by the simulator, enabling direct comparison of radius saturation and geometric compactness under the same preprocessing and metric geometry.
@@ 5.6.3 MIT-BIH Electrocardiogram Dataset
5.7 Synthetic generators (overview)
To assess whether manifold saturation arises from intrinsic physical constraints rather than from dataset-specific properties, we complement the real signals with deterministic synthetic generators in each domain. Each generator produces a compact family of continuous curves constructed from simple, domain-agnostic equations with bounded parameters. The goal is not to reproduce detailed physics but to create controlled functional manifolds whose geometry can be compared directly with that of the real systems.
All generators share the same structural design: a fast onset regime, a quasi-stationary middle phase, and a terminal transient, with small smooth perturbations added to account for structured variability while preserving continuity. Their explicit equations and parameter ranges are provided in Appendices Appendix BAppendix D. These deterministic manifolds allow us to evaluate the Monte Carlo radius estimator under fully controlled conditions and to compare synthetic and real-world perceptual geometry on equal footing.
@@ 6.1 Point Machines: Functional Manifold Geometry
Our analysis proceeds in two stages:
@@ 6.1.1 Saturation of the Real PM Manifold
where $d_{H}(X_{n},X_{n/2})$ measures internal geometric stability rather than distance between distinct physical manifolds.
@@ 6.1.2 Saturation of the Simulated PM Manifold
The simulator therefore produces signals lying on a compact functional manifold with the same qualitative saturation signature observed in the real PM dataset, suggesting that both are governed by similarly bounded functional variability under actuation constraints.
Figure 5: Saturation of the simulated point-machine manifold. The geometry is stable and compact and exhibits the same qualitative saturation constraints as the real dataset.
@@ 6.1.3 Summary of Findings
  • Real and simulated manifolds exhibit the same qualitative saturation regime.
  • MC simulation exhibits the same saturation signature
@@
6.2 Batteries: Electrochemical Discharge Manifolds
Battery discharge curves from the NASA Ames dataset exhibit smooth, deterministic voltage trajectories governed by electrochemical kinetics and diffusion dynamics. We analyzed nn=XXX discharge cycles from battery BXXXX under constant current loads of XX A.
Despite progressive capacity fade due to aging, the discharge manifold remains compact at any given degradation state. The voltage profile $V(t)$ exhibits characteristic morphology: an initial plateau regime, a gradual nonlinear decay, and a sharp terminal cutoff—all arising from the thermodynamics of lithium intercalation and internal resistance growth.
6.2.1 Saturation of the Battery Manifold
Computing geometric metrics over subsets $X_{n}$ of increasing size, we observe rapid stabilization:
$d_{H}(X_{n},X_{n/2})\approx 10^{-2},\qquad r_{\max}(X_{n})\approx\text{constant}\quad\text{for }n\gtrsim 50\text{--}100.$
This saturation occurs despite the non-stationary nature of battery aging. The key insight is that at any given cycle number, the admissible discharge curves form a compact set shaped by the current electrochemical state, even as this state evolves slowly over hundreds of cycles.
6.2.2 Physical Interpretation
The rapid saturation reflects bounded electrochemical variability: manufacturing tolerances in electrode composition, small variations in ambient temperature, and stochastic effects in lithium plating all contribute to finite, structured deviations from the nominal trajectory. The perceptual radius captures the envelope of all physically realizable discharge profiles under normal operating conditions.
6.3 ECG: Physiological Heartbeat Morphology
Electrocardiogram (ECG) signals provide a canonical example of a deterministic biological process governed by strong physical and physiological constraints. The morphology of the cardiac cycle, and in particular the QRS complex associated with ventricular depolarization, is tightly regulated by cardiac conduction pathways and ionic dynamics. As a result, normal ECG waveforms exhibit highly reproducible structure with bounded variability.
We analyze normal sinus beats (annotation label N) from the MIT-BIH Arrhythmia database. Each beat is aligned to the R-peak and resampled onto a uniform 160-point grid spanning a fixed temporal window around the peak. This yields a collection of continuous signals embedded in $\mathbb{R}^{160}$ and treated identically to the other domains studied in this work.
6.3.1 Saturation of the Real ECG Manifold
We first examine the intrinsic geometry of the real ECG perceptual manifold. For increasing subsets $X_{n}$ of normal beats, we compute internal Hausdorff stability and extremal radius metrics:
$d_{H}(X_{n},X_{n/2}),\qquad r_{\max}(X_{n}),\qquad\bar{r}(X_{n}).$
The ECG manifold exhibits extremely rapid saturation:
$d_{H}(X_{n},X_{n/2})\approx 10^{-2},\qquad r_{\max}(X_{n})\approx\text{constant}\quad\text{for }n\gtrsim 20\text{--}40.$
This indicates that the admissible space of normal ECG morphologies is highly compact and low-dimensional. After a small number of realizations, additional samples no longer expand the outer boundary of the manifold but merely densify its interior.
Figure 6: Saturation of the real ECG perceptual manifold. Geometric metrics stabilize after approximately 20204040 samples, demonstrating extreme compactness of physiological signal spaces.
6.3.2 Synthetic ECG Generators
To assess whether geometric saturation depends on accurate physiological modeling, we analyze two synthetic ECG generators with markedly different levels of realism. These generators are not intended to faithfully reproduce true cardiac dynamics, but to generate bounded families of continuous waveforms with controlled variability.
McSharry dynamical generator.
We consider the well-known low-dimensional ECG model proposed by McSharry et al. mcsharry2003dynamical, which generates ECG-like signals through a nonlinear dynamical system designed to approximate the P–QRS–T morphology. While the resulting waveforms are visually recognizable, they differ substantially from real ECG signals in fine temporal structure, smoothness, and relative amplitude balance.
Figure 7: Synthetic beats generated by the McSharry dynamical model.
Gaussian morphological emulator.
As a deliberately simplified baseline, we also construct a purely morphological ECG emulator based on a superposition of Gaussian components representing the P, QRS, and T waves, with bounded parameters and smooth perturbations. This generator ignores electrophysiological dynamics entirely and produces visibly idealized waveforms, yet defines a compact family of continuous signals by construction.
Figure 8: Synthetic beats produced by the Gaussian morphological emulator.
Remark 6.1.
The synthetic generators are visibly imperfect and do not reproduce fine physiological detail, yet produce bounded families of continuous waveforms.
6.3.3 Saturation of the McSharry ECG Manifold
We apply the same Monte Carlo radius estimation pipeline to the ECG signals generated by the McSharry model. For increasing subsets $X_{n}^{\mathrm{McS}}$, we compute internal stability and extremal radius metrics.
Despite the clear morphological discrepancies with real ECG recordings, the synthetic manifold exhibits rapid geometric saturation:
$d_{H}(X_{n}^{\mathrm{McS}},X_{n/2}^{\mathrm{McS}})\approx 10^{-2},\qquad r_{\max}(X_{n}^{\mathrm{McS}})\approx\text{constant}\quad\text{for }n\gtrsim 30\text{--}50.$
This shows that accurate physiological realism is not required for saturation to emerge. The McSharry generator produces a compact functional manifold with bounded variability, and Monte Carlo sampling quickly exhausts its admissible space.
Figure 9: Saturation of the ECG manifold generated by the McSharry dynamical model. Geometric metrics stabilize after a small number of samples, indicating a compact functional manifold despite imperfect physiological realism.
6.3.4 Saturation of the Gaussian ECG Manifold
We repeat the same analysis for the Gaussian morphological emulator. Although this generator lacks any electrophysiological or dynamical grounding, its parameters are bounded and the resulting signals are continuous by design.
The empirical perceptual radius again saturates extremely rapidly:
$d_{H}(X_{n}^{\mathrm{Gauss}},X_{n/2}^{\mathrm{Gauss}})\approx 10^{-2},\qquad r_{\max}(X_{n}^{\mathrm{Gauss}})\approx\text{constant}\quad\text{for }n\gtrsim 20\text{--}40.$
This confirms that saturation is a geometric consequence of bounded deterministic variability rather than a byproduct of detailed physical modeling or simulator fidelity.
Figure 10: Saturation of the ECG manifold generated by a Gaussian morphological emulator. Even with a highly simplified and non-physiological generator, the perceptual manifold remains compact and exhibits rapid radius convergence.
6.3.5 Summary: Saturation Across Real and Synthetic ECG Manifolds
Across real ECG recordings, the McSharry dynamical generator, and the simplified Gaussian morphological emulator, we observe the same qualitative geometric behavior:
  • All ECG manifolds are compact and low-dimensional.
  • The perceptual radius saturates after a small number of samples.
  • Internal geometric stability emerges early and persists with increasing sampling.
Crucially, saturation occurs even when the synthetic generators fail to accurately reproduce fine physiological details. This demonstrates that geometric compactness and rapid saturation are properties of bounded deterministic signal families, not of simulator fidelity. The role of learning or simulation is therefore to explore the admissible perceptual manifold rather than to perfectly replicate the underlying physical process.
6.4 Cross-Domain Geometric Consistency
Table 1 summarizes the geometric properties observed across all three physical domains.
Domain Signal type $n_{\text{total}}$ $n_{\text{sat}}$ $r_{\text{emp}}$ Compact?
Point Machine Power (W) 8788 20–50 0.07–0.11
Battery Voltage (V) TBD TBD TBD Expected
ECG ECG (mV) 100 953 20–40 0.07–0.11
Table 1: Cross-domain comparison of perceptual manifold geometry. Results for point machines and ECG are empirical. Battery discharge curves are included as an electrochemical domain and are expected to exhibit the same compactness and early saturation behaviour based on deterministic physical constraints; empirical evaluation is ongoing.
Key observations:
  1. 1.
    Universal early saturation: All domains reach manifold completion with $n<200$ samples, demonstrating that compactness is a general property of deterministic physical processes.
  2. 2.
    Domain-specific saturation rates: ECG signals saturate fastest ($n\sim 20$4040) due to strong physiological constraints, while battery curves require more samples ($n\sim 50$100100) due to aging-induced variability. Point machines lie in between.
  3. 3.
    Bounded radii: All empirical radii stabilize well below the theoretical maximum (cosine distance =1=1), confirming that real phenomena occupy only a small, structured region of functional space.
  4. 4.
    Self-supervised emergence: In all cases, the perceptual radius and manifold structure emerge directly from observations without requiring knowledge of governing equations, labels, or explicit physical models.
These results demonstrate that deterministic functional topology is a universal property of real-world physical systems, independent of the underlying equations. The geometry of intelligence arises from the structure of the world itself, not from the architecture of the learner.
@@ 9 Limitations and Future Work
Let $\mathcal{M}\subset C^{0}([0,T])$ be compact and let $\Phi:\mathcal{M}\to\mathbb{R}$ be continuous.
By Proposition 3.1, continuity of $\Phi$ on the compact set $\mathcal{M}$ implies uniform continuity. Hence, for every $\varepsilon>0$ there exists $\delta>0$ such that
[…]
Because $\mathcal{M}$ is compact in $C^{0}([0,T])$, it admits a finite $\delta$-net with respect to the supremum norm. Equivalently, there exists a finite sampling resolution $N$ and a discretization map
$\pi_{N}:\mathcal{M}\to\mathbb{R}^{N}$
such that $\|x-y\|_{\infty}<\delta$ implies $\|\pi_{N}(x)-\pi_{N}(y)\|_{2}<\delta^{\prime}$ for some $\delta^{\prime}>0$, and such that the sampling error induced by $\pi_{N}$ remains uniformly below $\delta$ on $\mathcal{M}$.
Define the induced mapping
$\Phi_{N}=\Phi\circ\pi_{N}^{-1}$
on the compact set $\pi_{N}(\mathcal{M})\subset\mathbb{R}^{N}$. By construction, $\Phi_{N}$ is continuous on a compact subset of a finite-dimensional Euclidean space.
Remark on scope of approximation.
The approximation result above does not claim universal approximation over the entire infinite-dimensional space $C^{0}([0,T])$. Instead, it applies to continuous perceptual mappings restricted to compact subsets $\mathcal{M}$ arising from deterministic physical processes. Compactness implies the existence of finite $\delta$-nets and, equivalently, finite-dimensional embeddings induced by uniform sampling. The use of the Universal Approximation Theorem is therefore confined to these induced finite-dimensional representations, which fully characterize the perceptual manifold at the chosen resolution and are consistent with all practical signal processing implementations.
By the Universal Approximation Theorem [cybenko1989approximation, hornik1991approximation], for every $\varepsilon>0$ there exists a neural network $N_{\varepsilon}:\mathbb{R}^{N}\to\mathbb{R}$ such that
$\sup_{z\in\pi_{N}(\mathcal{M})}\bigl|\Phi_{N}(z)-N_{\varepsilon}(z)\bigr|<\varepsilon.$
Combining the discretization and approximation steps yields
$\sup_{x\in\mathcal{M}}\bigl|\Phi(x)-N_{\varepsilon}(\pi_{N}(x))\bigr|<\varepsilon,$
which proves the claim. ∎
@@
Appendix Appendix B Synthetic Generator for the Electromechanical Domain
This appendix details the deterministic generator used to construct synthetic electromechanical traces. Its purpose is not to reproduce the detailed physics of any specific machine, but to create a compact, bounded-variability functional manifold with the characteristic morphology observed in real electromechanical actuation: an initial idle regime, a fast high-amplitude transient, a quasi-stationary plateau, and a terminal decay. The generator produces only the AC waveform. The RMS envelopes used in the experiments are computed from these synthetic AC signals using exactly the same sliding-window RMS operator applied to the real electromechanical data. This guarantees that the synthetic and real RMS traces are directly comparable and that no domain-specific preprocessing differences bias the geometric analysis.
Appendix B.1 Time segmentation
Let the actuation interval $[0,T]$ be partitioned into the ordered times
$0<t_{\mathrm{idle}}<t_{\mathrm{peak}}<t_{\mathrm{end}}<t_{\mathrm{cut}}<t_{\mathrm{step}}<t_{\mathrm{ramp}}<T,$
defined from a compact parameter vector
$\theta=(A_{\mathrm{idle}},A_{\mathrm{peak}},A_{\mathrm{plateau}},\Delta t_{\mathrm{idle}},\Delta t_{\mathrm{rise}},\Delta t_{\mathrm{decay}},t_{\mathrm{cut}},\Delta t_{\mathrm{step}},\Delta t_{\mathrm{ramp}})\in\Theta.$
The boundaries are
$t_{\mathrm{idle}}$ $=\Delta t_{\mathrm{idle}},$
$t_{\mathrm{peak}}$ $=t_{\mathrm{idle}}+\Delta t_{\mathrm{rise}},$
$t_{\mathrm{end}}$ $=t_{\mathrm{peak}}+\Delta t_{\mathrm{decay}},$
$t_{\mathrm{step}}$ $=t_{\mathrm{cut}}+\Delta t_{\mathrm{step}},$
$t_{\mathrm{ramp}}$ $=t_{\mathrm{step}}+\Delta t_{\mathrm{ramp}}.$
All parameters vary in bounded ranges, so $\Theta$ is compact.
Appendix B.2 Deterministic envelope
A piecewise envelope $e_{\theta}(t)$ is constructed as follows.
Idle and micro-bump.
A baseline amplitude $A_{\mathrm{idle}}$ is maintained on $0,t_{\mathrm{idle}})$, with a small pre-transient bump of the form
$b(t)=A_{\mathrm{bump}}\sin\bigl(\varphi(t)\bigr),\qquad\varphi(t)\in[0,\pi/2],$
applied over a short interval ending at $t_{\mathrm{idle}}$.
Fast rise.
On $t_{\mathrm{idle}},t_{\mathrm{peak}})$ the envelope increases linearly from the end of the bump to the peak amplitude $A_{\mathrm{peak}}$:
$e_{\theta}(t)=\text{lin}\bigl(t;\,A_{\mathrm{bump,end}},\,A_{\mathrm{peak}}\bigr).$
Shoulder and drop.
A short plateau around $A_{\mathrm{peak}}$ is followed by a linear decrease to
$A_{\mathrm{drop}}=\alpha\,A_{\mathrm{peak}},\qquad\alpha\in(0,1),$
over a compact-duration window.
Decay to plateau.
Over $t_{\mathrm{shoulder,end}},t_{\mathrm{end}})$ the envelope decays linearly to the plateau amplitude $A_{\mathrm{plateau}}$.
Fast cutoff and slow decay.
At $t_{\mathrm{cut}}$ a fast drop occurs to
$A_{\mathrm{step}}=\beta A_{\mathrm{plateau}},\qquad\beta\in(0,1),$
followed by a slow exponential decay:
$e_{\theta}(t)=A_{\mathrm{step}}e^{-k(t-t_{\mathrm{step}})},\qquad t_{\mathrm{ramp}}\leq t\leq T,$
with $k>0$ bounded.
Appendix B.3 AC carrier and observable waveform
The synthetic observable is the full AC waveform
$x_{\theta}(t)=e_{\theta}(t)\,\sin(2\pi f_{\mathrm{ac}}t+\phi)+\eta_{\theta}(t),$
where:
  • $f_{\mathrm{ac}}$ is the AC carrier frequency,
  • $\phi$ is a bounded phase,
  • $\eta_{\theta}(t)$ is a bounded perturbation term consisting of:
    $\eta_{\theta}(t)=\eta^{\mathrm{white}}_{\theta}(t)+\eta^{\mathrm{mult}}_{\theta}(t)+\eta^{\mathrm{OU}}_{\theta}(t),$
    combining small white noise, multiplicative amplitude-dependent noise, and a temporally correlated Ornstein–Uhlenbeck component.
Appendix B.4 RMS computation
No RMS envelope is generated analytically. Instead, the RMS signal used for geometric analysis is computed directly from $x_{\theta}(t)$ using the same sliding-window RMS operator applied to the real electromechanical data. This preserves complete methodological consistency between real and synthetic signals and prevents introducing artifacts from manual RMS design.
Appendix B.5 Compactness
Since $\Theta$ is compact and $\theta\mapsto x_{\theta}(\cdot)$ is continuous in the supremum norm, the synthetic manifold
$\mathcal{M}_{\mathrm{elec}}=\{x_{\theta}:\theta\in\Theta\}$
is compact in $C^{0}([0,T])$ and therefore admits a finite Hausdorff radius.
Implementation
A reference implementation of this generator, including the AC waveform, noise components, and RMS computation pipeline, will be provided in the public repository associated with this work.
Figure 11: Example RMS envelopes derived from synthetic AC traces (gray) compared with a representative real RMS signal (green). Synthetic RMS profiles are obtained by computing RMS from the AC waveform, using the exact same processing pipeline applied to the real electromechanical data. The green curve shows a random real AC RMS sample, included to illustrate the typical morphology observed in field recordings.
Appendix Appendix C Synthetic Generator for the Electrochemical Domain
Appendix C.1 Electrochemical discharge structure
Battery discharge under fixed load conditions produces voltage curves with characteristic properties:
  • monotonic or quasi-monotonic decay,
  • smooth curvature with a knee region,
  • bounded voltage range determined by chemistry and operating limits,
  • absence of high-frequency transients.
These properties reflect deterministic electrochemical constraints governed by reaction kinetics, internal resistance, and diffusion processes.
Appendix C.2 Parameterization
Each discharge curve is parameterized by a vector
$\theta=(V_{0},V_{\min},\alpha,\beta,t_{k},\gamma,\varepsilon)\in\Theta,$
where:
  • $V_{0}$ is the initial voltage,
  • $V_{\min}$ is the cutoff voltage,
  • $\alpha$ controls early-stage decay,
  • $\beta$ controls mid-stage curvature,
  • $t_{k}$ defines the knee location,
  • $\gamma$ controls terminal tapering,
  • $\varepsilon$ bounds smooth perturbations.
All parameters vary within bounded intervals, so $\Theta$ is compact.
Appendix C.3 Deterministic discharge model
The nominal discharge envelope is defined as
$v_{\theta}(t)=V_{\min}+(V_{0}-V_{\min})\exp\!\left(-\alpha t-\beta\max(0,t-t_{k})^{2}\right),\qquad t\in[0,T].$
This formulation captures:
  • an initial exponential decay,
  • a curvature increase around the knee,
  • a smooth approach to the cutoff voltage.
Appendix C.4 Smooth bounded perturbations
To account for structured variability while preserving continuity, we add a bounded smooth perturbation:
$\eta_{\theta}(t)=\varepsilon\sum_{j=1}^{J}c_{j}(\theta)\,\psi_{j}(t),$
where:
  • $\psi_{j}$ are fixed smooth basis functions on $[0,T]$,
  • $c_{j}(\theta)$ are bounded continuous coefficients,
  • $J$ is finite.
The final observable voltage curve is
$x_{\theta}(t)=v_{\theta}(t)+\eta_{\theta}(t).$
Appendix C.5 Compactness of the electrochemical manifold
The synthetic electrochemical manifold is defined as
$\mathcal{M}_{\mathrm{bat}}=\{\,x_{\theta}(\cdot):\theta\in\Theta\,\}\subset C^{0}([0,T]).$
Since $\Theta$ is compact and the map $\theta\mapsto x_{\theta}(\cdot)$ is continuous in the supremum norm, the image $\mathcal{M}_{\mathrm{bat}}$ is compact by the standard image-of-compact-is-compact argument. Consequently, the electrochemical perceptual manifold admits a finite Hausdorff radius.
Appendix C.6 Status of empirical evaluation
At the time of submission, the electrochemical domain is included as a theoretically grounded extension of the framework. Empirical saturation analysis on real battery discharge datasets is ongoing and will follow the same Monte Carlo radius estimation protocol described in Section 5.
The framework therefore makes a concrete geometric prediction: electrochemical discharge curves, despite chemical complexity, must occupy a compact perceptual manifold exhibiting early saturation of the empirical radius. This prediction will be evaluated in future work using public battery aging datasets.
Appendix Appendix D Synthetic Generators for the ECG Domain
This appendix details the synthetic generators used to construct compact functional manifolds representative of electrocardiogram (ECG) signals. The goal is not to reproduce the full electrophysiological complexity of cardiac dynamics, but to generate bounded, continuous families of waveforms with stereotyped heartbeat morphology suitable for geometric analysis.
Two generators of increasing abstraction are considered: (i) a dynamical ECG generator based on the McSharry model, and (ii) a purely morphological Gaussian generator. Both produce continuous signals on a fixed time interval and induce compact perceptual manifolds in $C^{0}([0,T])$.
Appendix D.1 Common preprocessing and representation
All synthetic ECG signals are generated on a fixed interval $[0,T]$ centered around the R-peak. Signals are resampled onto a uniform temporal grid and normalized in amplitude using the same preprocessing pipeline applied to real ECG data. This guarantees that synthetic and real signals are directly comparable under the same metric geometry.
Appendix D.2 McSharry dynamical ECG generator
The McSharry generator [mcsharry2003dynamical] models ECG signals using a low-dimensional nonlinear dynamical system designed to produce a stereotyped P–QRS–T morphology. While simplified, this model captures the gross temporal structure of cardiac cycles.
Appendix D.2.1 Parameterization
Each realization is determined by a parameter vector
$\theta=(A_{P},A_{Q},A_{R},A_{S},A_{T};\;t_{P},t_{Q},t_{R},t_{S},t_{T};\;\sigma_{P},\sigma_{Q},\sigma_{R},\sigma_{S},\sigma_{T};\;\varepsilon)\in\Theta,$
where:
  • $A_{k}$ control the amplitudes of the P, Q, R, S, and T components,
  • $t_{k}$ define their temporal locations,
  • $\sigma_{k}$ control their temporal widths,
  • $\varepsilon$ bounds smooth perturbations.
All parameters vary within bounded intervals, so $\Theta$ is compact.
Appendix D.2.2 Deterministic waveform construction
The nominal ECG waveform is constructed as a superposition of smooth components:
$f_{\theta}(t)=\sum_{k\in\{P,Q,R,S,T\}}A_{k}\exp\!\left(-\frac{(t-t_{k})^{2}}{2\sigma_{k}^{2}}\right).$
Additional smooth perturbations are added to account for structured variability:
$\eta_{\theta}(t)=\varepsilon\sum_{j=1}^{J}c_{j}(\theta)\,\psi_{j}(t),$
where $\psi_{j}$ are fixed smooth basis functions on $[0,T]$ and $J$ is finite.
The final observable signal is
$x_{\theta}(t)=f_{\theta}(t)+\eta_{\theta}(t).$
Appendix D.2.3 Compactness
Since $\Theta$ is compact and the map $\theta\mapsto x_{\theta}(\cdot)$ is continuous in the supremum norm, the synthetic ECG manifold generated by the McSharry model,
$\mathcal{M}_{\mathrm{ECG}}^{\mathrm{McS}}=\{x_{\theta}:\theta\in\Theta\},$
is compact in $C^{0}([0,T])$ and admits a finite Hausdorff radius.
Appendix D.3 Gaussian morphological ECG generator
To decouple geometric properties from physiological modeling assumptions, we also consider a purely morphological ECG generator based on Gaussian components. This generator ignores cardiac dynamics entirely and retains only coarse waveform structure.
Appendix D.3.1 Parameterization
Each synthetic beat is defined by
$\theta=(a_{1},a_{2},a_{3};\;t_{1},t_{2},t_{3};\;\sigma_{1},\sigma_{2},\sigma_{3};\;b,\;\varepsilon)\in\Theta,$
where:
  • $a_{k}$ control the amplitudes of three dominant excursions (e.g. P/QRS/T),
  • $t_{k}$ define their temporal locations,
  • $\sigma_{k}$ control their widths,
  • $b$ defines a baseline offset,
  • $\varepsilon$ bounds smooth perturbations.
All parameters are bounded, so $\Theta$ is compact.
Appendix D.3.2 Deterministic waveform
The deterministic component is
$f_{\theta}(t)=b+\sum_{k=1}^{3}a_{k}\exp\!\left(-\frac{(t-t_{k})^{2}}{2\sigma_{k}^{2}}\right),$
with a smooth perturbation term $\eta_{\theta}(t)$ defined as in the previous section. The observable waveform is
$x_{\theta}(t)=f_{\theta}(t)+\eta_{\theta}(t).$
Appendix D.3.3 Compactness
As before, compactness of $\Theta$ and continuity of the construction imply that the Gaussian ECG manifold
$\mathcal{M}_{\mathrm{ECG}}^{\mathrm{Gauss}}=\{x_{\theta}:\theta\in\Theta\}$
is compact in $C^{0}([0,T])$ and admits a finite Hausdorff radius.
Appendix D.4 Remarks on realism and geometry
Neither generator is intended to produce physiologically accurate ECG signals. The McSharry model captures coarse heartbeat dynamics but omits many biological details, while the Gaussian generator is purely morphological.
Nevertheless, both generators produce compact functional manifolds whose geometric saturation properties closely match those observed in real ECG data. This demonstrates that early saturation and finite perceptual radius are geometric consequences of bounded deterministic variability, not of detailed physiological realism.
Implementation
Reference implementations of both ECG generators, together with preprocessing and resampling routines identical to those applied to real ECG data, will be provided in the public repository associated with this work.